Obtain a four-term Taylor polynomial approximation valid near x=,0 for each (1+x)^-1/3. Estimate the range of x over which three term polynomials will give two decimal accuracy


1
Expert's answer
2021-04-26T17:47:11-0400

set "f(x)=(1+x)^{-1\/3}"

for a four term Taylor polynomial centered at x=0

"f(x)= (1+x)^{-1/3}\implies f(0)=1"

"f'(x)=-\frac 13 (1+x)^{-4/3}\implies f'(0)=-\frac 12"

"f''(x)= -\frac 49(1+x)^{-7/3} \implies f''(0)= \frac 49"

"f'''(x)= -\frac {28}{27}(1+x)^{-10/3}\implies f'''(0)= -\frac {28}{27}"

now the four term Taylor polynomial approximation to f(x) valid near x=0 is;

"f(x) =(1+x)^{-1/3} = 1-\frac 13x + \frac{\frac 49}{2!}x^2- \frac {\frac {28}{27}}{3!}x^3"

"= 1-\frac 13 x + \frac 29 x^2 - \frac {14}{81}x^3"

for the range of values of x over which these term polynomials will give a two decimal accuracy we may use the Lagrange error bound;

"m\frac {|x-a|^{n+1}}{(n+a)!}<=" Β maximum admissible error

where "m= max _{x\isin\Iota}|f^{n+1}(x)|"

"\therefore m \frac {|x-0|^3}{3!}<= 0.01" or "|x|^3<= \frac {0.01*3!}{m}"

"|x|<= \frac {0.39}{\sqrt [3]m}"

"\implies -0.4<= x<= 0.4" is the range desired.